# Enhancing Data Privacy in Large Language Models through Private Association Editing

### v1

@@ Abstract
Davide Venditti*${}^{\textbf{*}1}$, Elena Sofia Ruzzetti*${}^{\textbf{*}1}$, Giancarlo A. Xompero2 Cristina Giannone2, Andrea Favalli2, Raniero Romagnoli2 Fabio Massimo Zanzotto1 1University of Rome Tor Vergata, Italy 2Almawave S.p.A., Via di Casal Boccone, 188-190 00137, Rome, IT davide.venditti@uniroma2.it, elena.sofia.ruzzetti@uniroma2.it fabio.massimo.zanzotto@uniroma2.it
@@ 1 Introduction
A massive pretraining phase seems to be the key to obtaining versatility and accuracy in a large number of tasks: Large language models (LLMs) are indeed able to perform accurately many tasks by capturing information from their training data. Even in zero-shot scenarios, LLMs serve as alternative sources of information Hou et al. (2024), perform translation tasks Mu et al. (2023), translate natural language requests into code Ranaldi et al. (2024), and are definitely capable of capturing world knowledge Petroni et al. (2019, 2020). The massive pretraining phase seems to be the key to obtaining versatility and accuracy in this large variety of tasks.
However, growing larger, training data for LLMs have become uncontrollable and may inadvertently contain some private personal information of unaware people. LLMs may potentially retain this sensitive information Carlini et al. (2021, 2023); Huang et al. (2022). This is a potential threat in privacy of unaware people. Indeed, by performing Training Data Extraction attacks, Carlini et al. (2021) showed that LLMs may verbatim generate strings containing sensitive information observed during training. Then, attackers may gain access to private information.
@@
Strategies to remove sensitive information from LLMs are needed and mandatory, as preserving privacy is a must. Yet, the straight-forward technique of remove-and-retrain is unfeasible as extremely expensive.
In this paper, we propose Private Association Editing (PAE) to remove memorized private information adjusting parameters of LLMs without re-training (see Fig. 1). Stemming from MEMIT Meng et al. (2023b) formulation to edit factual knowledge, we define PAE as a novel model-editing defense strategy based on the idea of breaking the association between personal information and the identity of the person to whom it belongs. We anonymize the private information directly in the model, replacing the original information with masked – but semantically equivalent – information. We experiment with GPT-J Wang and Komatsuzaki (2021) as it is an open-source model that contains documented private information. We perform Training Data Extraction attacks Huang et al. (2022) before and after our model-editing defenses and we show that our strategies are an efficient alternative to make a model more robust against the generation of private information while keeping constant its performance in generating texts.
@@ 2 Background
Large Language Models (LLMs) are prone to emit private information. Indeed, attacking LLMs to extract memorized private information is possible by using black-box access to language models. Training Data Extraction (TDE) is a technique to extract this private information Carlini et al. (2021). It consists of querying the target model to force it to produce its own training data. A textual training example is considered "extractable" if a specific prefix can be used to prompt the model to generate the exact training example from its training set. Carlini et al. (2021) found that GPT-2 often retains and reveals personal information such as Twitter handles, email addresses, and Universal Unique Identifiers (UUIDs). The memorization of training examples explains the success of these attacks: when LLMs are prompted with a prefix encountered during training, they often complete the prompt with the remaining part of the training sequence Carlini et al. (2023).
@@
Attacks may be particularly effective in open LLMs. Huang et al. (2022) demonstrated that conditioning a model with a prompt that is part of the training data can result in the leakage of personal identifiable information (PII), such as email addresses. They also showed that this method is more effective than creating entirely new, unseen prompts. Nasr et al. (2023) revealed that Carlini et al. (2021) method is even more effective than previously expected. By querying open-source models, they confirmed the success of the attack procedure using the training data solely for verification purposes. They conducted these attacks on open models like GPT-Neo Black et al. (2022) and Pythia Biderman et al. (2023), starting with prompts sourced from Wikipedia.
Even closed LLMs may reveal private information by using Training Data Extraction. Since these attacks require only black-box access to the model, closed models like GPT-3.5 and GPT-4 can be attacked. In fact, using the same prompts proposed by Huang et al. (2022), Wang et al. (2024) demonstrate that GPT-3.5 and GPT-4 can predict respectively around 5% and 4% of the email addresses accurately.
As personal information leakage from LLMs is a concrete possibility, model editing is a possible solution as opposed to an expensive remove-and-retrain strategy.
@@
Model editing in LLMs refers to the process of modifying specific aspects of a model’s behavior or knowledge without retraining it from scratch. This involves making targeted adjustments to the model’s parameters or responses to correct errors, update information, or adapt to new requirements. Mitchell et al. (2022) introduced a semi-parametric editing methodology, employing a retrieval-augmented counterfactual model, that effectively modulates neural network predictions over the SERAC dataset. Cao et al. (2021) proposed KNOWLEDGEEDITOR that efficiently and reliably edits factual knowledge within language models, ensuring consistency across various formulations of facts. Furthermore, Yao et al. (2023) introduced MEND on various datasets, demonstrating its ability to rapidly and effectively edit large-scale models’ behaviors without extensive retraining. Since these methods can modify factual information memorized in LLMs, our goal is to exploit them to erase private information inadvertently ingested during training.
@@
Similarly to the method defined in our paper, Patil et al. (2023) investigated model editing techniques to modify the information memorized in LLMs concluding that information cannot be erased. In particular, they applied TDE attacks against the GPT-J Wang and Komatsuzaki (2021) model and demonstrated that in black-box access–performing attacks that also include paraphrases of the original prompt– model editing cannot erase factual information memorized in GPT-J. Our setting is different: in fact, Patil et al. (2023) investigated the effectiveness of model editing only on factual information from sentences derived from Wikipedia, and not directly present in the training data – the Pile Gao et al. (2020). By definition, the model under attack does not verbatim memorize information that is not in training data: since the examples used by Patil et al. (2023) are derived from Wikipedia and not included in the Pile, while the factual information they contain is memorized, they cannot be verbatim memorized. In our experiments, we directly study the effectiveness of model editing to delete private information that is verbatim memorized with a focus on privacy rather than factual information.
To the best of our knowledge, this is a novel approach to protect private LLMs from personal information leakage that show the potential beneficial effect on privacy preserving.
@@ 3.1 Training Data Extraction Attacks to recover Sensitive Information
To detect the presence of memorized Personally Identifiable Information LLMs, we follow the attack pipeline and attack prompts defined by Huang et al. (2022). They defined two kinds of attacks depending on how information is stored and retrieved: (1) a model memorizes personal information if there exists a prompt from the training data that leads the model to generate that information; (2) in contrast, a model associates an individual to its personal information if there exists a prompt not seen during training but containing a reference to an individual that leads to the generation of PII. Huang et al. (2022) already demonstrated that memorization is more common in LLM than association, showing that a model from the GPT-Neo [^1] https://www.eleuther.ai/artifacts/gpt-neo family can predict emails more accurately when conditioned with prompts from the training data rather than wthen analyzely unseen prompts.
@@ 3.2 Private Association Editing as Efficient Defense against Privacy Attacks
Our PAE employs model editing techniques based on ROME (Rank-One Model Editing) Meng et al. (2023a) and MEMIT (Model Editing via Iterative Training) Meng et al. (2023b) as a defensive strategy against attacks aimed at safeguarding the sensitive data used to train Large Language Models (LLMs). Then, the scalability to editing different facts in a batch is facilitated by the MEMIT framework, which allows us to incorporate as many elements in the form of modifications as desired. These modifications are seamlessly executed on the same model also avoiding degradation of the model’s performance.
@@
The procedure to edit a private association uses PAE cards based on the MEMIT modification card. The basic structure of a MEMIT modification card is composed of a prompt, a ground truth, a target, and a subject. Our PAE cards specialize the MEMIT modification card on a particular PII. We have defined two main types of PAE cards to mask the private information of users. The first type is called "explicit" because it directly identifies the connection between the person and the private data and perfectly adheres to the MEMIT implementation. For example, an explicit prompt is "{name} has an email address that is". The second type is "implicit" which features a prompt that does not necessarily include the person’s name as the subject of the sentence, favoring a more precise meaning of the sentence. An example of an implicit prompt is "The mail address of {name} is".
@@
It is important to note that we used MEMIT in "batch" mode because we are interested in fixing the model and subjecting it to $k$ modifications. In this way, we are able to use MEMIT performing $k$ modifications at the same time, instead of performing single edits separately and recreating the model based on the post-edited weights obtained from the last edit every time.
In a real-world scenario, recreating or retraining a model for each requested modification is not feasible. Instead, with our strategy called "one model, n edits" we are able to make all requested changes to a single model. By masking and anonymizing the email address, we make it more challenging for attackers to elicit specific private data from the model in response to particular prompts. This methodology effectively reduces the risk of sensitive information being inadvertently disclosed by the model.
@@ 3.3 Evaluating Language Modeling Performance
The final step of the procedure for preserving privacy with PAE is to investigate whether the LLM maintains its behavior in text generation. In fact, Model Editing techniques, in general, and PAE, in particular, may perturb the language model capabilities due to the intervention on the model parameters.
The LLM assessment procedure we describe in this Section aims to verify that the privacy-preserving language model is not a worse model than the original one. The main idea is that LLMs capabilities are not perturbed if people are not able to determine which of the two models is responsible for which generation, then it means that the edit procedure does not affect model performance: there is no one better than the other, a user of the system would be equally happy to use one or the other.
@@ 4.1 Experimental Setup
In this section, we discuss the parameters of our experiment to allow replicability: the analized LLM and related datasets, the intricacies of MEMIT used in our PEA, and, finally, the set-up of the evaluation of the LLMs.
@@
In our experiments, we test the GPT-J model Wang and Komatsuzaki (2021) that is designed to generate human-like text continuations from prompts: it is a large model, with 6 billion parameters. This model is trained on an open dataset, the Pile Gao et al. (2020). The Pile is a diverse, large-scale text corpus that aggregates various sources, including books, articles, websites, and scientific papers. It spans multiple languages and domains, making it an ideal training resource for language models like GPT-J. The Pile contains a rich variety of text, enabling the model to learn from a wide range of contexts and topics. One of the constituent sub-datasets within The Pile is the Enron Emails Klimt and Yang (2004) corpus. This dataset contains text from approximately 150 users, primarily senior management of Enron, organized into folders. It includes a total of about 0.5 million email messages. The Enron Emails dataset was originally made public during the investigation into Enron’s accounting methods. Its inclusion in the Pile mimic the inadvertently insertion into the training data of private information, in particular of PII like email adressess. For this reason, the Enron Email dataset represents a natural starting point to test GPT-J memorization of PII.
Intricacies of MEMIT
@@
There are two distinct ways to apply model editing using MEMITYao et al. (2023) given $N$ elements to modify: batch and sequential editing. Batch editing involves editing $k$ elements in an LLM simultaneously. Conversely, sequential editing focuses on editing $N$ elements within an LLM in a sequential way, with each edit on a subset of the $N$ elements, performed on the new model retaining previous edits. While batch editing may be sufficient to preserve privacy, the sequential editing approach is closer to the real-world need to constantly update model parameters, as more privacy leakages may be discovered over time.
@@
In our research, we initially adopt the batch editing approach with $k=N$. This approach is the safest – in principle – since the post-edited parameters are directly the pretrained ones. Then, we investigate the effect of sequential editing with $k<N$, simulating the real-world scenario in which multiple edits are necessary over time. For PAE to be applicable, in both scenario our method should lead to a comparable decrease in privacy leaks.
Evaluation of post-edited LLM
@@
For the automatic assessment procedure, we measure the difference in generations for the pre-trained GPT-J model and the post-edited version by generating a small paragraph starting from 45 prompts extracted from the Book3 Rae et al. (2022) dataset, included in the Pile . We prompted the post-edited models and the pretrained one with 20 tokens of the 45 randomly selected examples and we evaluate how similar the generations are measuring their overlap. he higher the similarity, the higher the likelihood that the PAE does not influence the overall performance of the model. Evaluation measures are the ROUGE and the METEOR scores.
@@
For the manual assessment procedure, we generate with post-edited models and with the pretrained one a short paragraph from 10 different prompts (a complete list can be found in the Appendix 6.1). We collect the generations for the pre-edit model and the model post-edited according to each of the editing strategies. Hence, in total, we collect 30 generations. Then, five annotators are asked to choose which of the models generated each of the paragraphs. Three sample generations of each model were provided, and the annotators were informed that two out of three models had been post-edited, but none of them were informed which of the three systems had been post-edited. Evaluation measures are the classification accuracy of each annotator and the Fleiss’ K inter-annotator agreement: a low score on both can confirm that the models are indistinguishable.

### v2

@@ Abstract
Davide Venditti*${}^{\textbf{*}1}$, Elena Sofia Ruzzetti*${}^{\textbf{*}1}$, Giancarlo A. Xompero1,2 Cristina Giannone2, Andrea Favalli2, Raniero Romagnoli2 Fabio Massimo Zanzotto1 1University of Rome Tor Vergata, Italy 2Almawave S.p.A., Via di Casal Boccone, 188-190 00137, Rome, IT davide.venditti@uniroma2.it, elena.sofia.ruzzetti@uniroma2.it fabio.massimo.zanzotto@uniroma2.it
@@ 1 Introduction
A massive pretraining phase seems to be the key to obtaining versatility and accuracy in a large number of tasks: Large language models (LLMs) are indeed able to perform accurately many tasks by capturing information from their training data. Even in zero-shot scenarios, LLMs serve as alternative sources of information Hou et al. (2024), perform translation tasks Mu et al. (2023), translate natural language requests into code Ranaldi et al. (2024), and are definitely capable of capturing world knowledge Petroni et al. (2019, 2020). The massive pretraining phase seems to be the key to obtaining versatility and accuracy in this large variety of tasks. However, growing larger, training data for LLMs have become uncontrollable and may inadvertently contain some private personal information of unaware people. LLMs may potentially retain this sensitive information Carlini et al. (2021, 2023); Huang et al. (2022). This is a potential threat to the privacy of unaware people. Indeed, by performing Training Data Extraction attacks, Carlini et al. (2021) showed that LLMs may verbatim generate strings containing sensitive information observed during training. Then, attackers may gain access to private information.
@@
Strategies to remove sensitive information from LLMs are needed and mandatory, as preserving privacy is a must. Yet, the straightforward technique of remove-and-retrain is extremely expensive.
In this paper, we propose Private Association Editing (PAE): a novel “one model, $k$ edits” strategy to remove memorized private information adjusting parameters of LLMs without re-training (see Fig. 1). Stemming from MEMIT Meng et al. (2023b) formulation to edit factual knowledge, we define PAE as a novel model-editing defense strategy based on the idea of breaking the association between personal information and the identity of the person to whom it belongs. We anonymize the private information directly in the model, replacing the original information with masked – but semantically equivalent – information. We experiment with GPT-J Wang and Komatsuzaki (2021) as it is an open-source model that contains documented private information. We perform Training Data Extraction attacks Huang et al. (2022) before and after our model-editing defenses, and we show that our strategies are an efficient alternative to make a model more robust against the generation of private information while keeping constant its performance in generating texts. In particular, we empirically demonstrate the feasibility of our “one model, $k$ edits” approach: we aim to protect the privacy of multiple users with a single edit, successfully applying model editing both in batch and in sequence.
@@ 2 Background
Large Language Models (LLMs) are prone to emit private information. Indeed, attacking LLMs to extract memorized private information is possible by using black-box access to language models. Training Data Extraction (TDE) is a technique to extract this private information Carlini et al. (2021). It consists of querying the target model to force it to produce its own training data. A textual training example is considered "extractable" if a specific prefix can be used to prompt the model to generate the exact training example from its training set. Carlini et al. (2021) found that GPT-2 often retains and reveals personal information such as Twitter handles, email addresses, and Universal Unique Identifiers.The memorization of training examples explains the success of these attacks: when LLMs are prompted with a prefix encountered during training, they often complete the prompt with the remaining part of the training sequence Carlini et al. (2023).
@@
Attacks may be particularly effective in open LLMs. Huang et al. (2022) demonstrated that conditioning a model with a prompt that is part of the training data can result in the leakage of personally identifiable information (PII), such as email addresses. They also showed that this method is more effective than creating entirely new, unseen prompts. Nasr et al. (2023) revealed that Carlini et al. (2021) method is even more effective than previously expected: by querying open-source models like GPT-Neo Black et al. (2022) and Pythia Biderman et al. (2023), they confirmed the success of the attack procedure using the training data solely for verification purposes. Since these attacks require only black-box access to the model, closed models like GPT-3.5 and GPT-4 can be successfully attacked Wang et al. (2024).
As personal information leakage from LLMs is a concrete possibility, different strategies have been explored to avoid a model generating potentially harmful content. Yao et al. (2024) propose an unlearning mechanism that requires only negative samples – i.e., examples in which the model generates harmful content – to stop the generation of undesirable outputs. However, as the majority of machine unlearning approaches Liu et al. (2024), it requires the definition of a retain set that contains samples used to preserve the utility of the model. Our aim is to modify only a batch of information without further training or additional data.
@@
Model editing is a possible solution as opposed to an expensive remove-and-retrain strategy or unlearning strategies. Model editing in LLMs refers to the process of modifying specific aspects of a model’s behavior or knowledge without retraining it from scratch. This involves making targeted adjustments to the model’s parameters or responses to correct errors, update information, or adapt to new requirements. Mitchell et al. (2022) introduced a semi-parametric editing methodology, employing a retrieval-augmented counterfactual model. Cao et al. (2021) proposed KNOWLEDGEEDITOR that efficiently and reliably edits factual knowledge within language models, ensuring consistency across various formulations of facts. Furthermore, Yao et al. (2023) introduced MEND on various datasets, demonstrating its ability to rapidly and effectively edit large-scale models’ behaviors without extensive retraining. Since these methods can modify factual information memorized in LLMs, our goal is to exploit them to erase private information inadvertently ingested during training.
@@
Similarly to the method defined in our paper, Patil et al. (2023) investigated model editing techniques to modify the information memorized in LLMs, concluding that information cannot be erased. In particular, they applied TDE attacks against the GPT-J Wang and Komatsuzaki (2021) model and demonstrated that in black-box access–performing attacks that also include paraphrases of the original prompt– model editing cannot erase factual information memorized in GPT-J. Our setting is different: in fact, Patil et al. (2023) investigated the effectiveness of model editing only on factual information from sentences derived from Wikipedia and not directly present in the training data – the Pile Gao et al. (2020). By definition, the model under attack does not verbatim memorize information that is not in training data: since the examples used by Patil et al. (2023) are derived from Wikipedia and not included in the Pile, while the factual information they contain is memorized, they cannot be verbatim memorized. In our experiments, we directly study the effectiveness of model editing in deleting private information that is verbatim memorized with a focus on privacy rather than factual information.
@@ 3.1 Training Data Extraction Attacks to recover Sensitive Information
To detect the presence of memorized Personally Identifiable Information LLMs, we follow the attack pipeline and attack prompts defined by Huang et al. (2022). They defined two kinds of attacks depending on how information is stored and retrieved: (1) a model memorizes personal information if there exists a prompt from the training data that leads the model to generate that information; (2) in contrast, a model associates an individual to its personal information if there exists a prompt not seen during training but containing a reference to an individual that leads to the generation of PII. Huang et al. (2022) already demonstrated that memorization is more common in LLM than association, showing that a model from the GPT-Neo [^1] https://www.eleuther.ai/artifacts/gpt-neo family can predict emails more accurately when conditioned with prompts from the training data rather than with unseen prompts.
@@ 3.2 Private Association Editing as Efficient Defense against Privacy Attacks
Our PAE employs model editing techniques based on MEMIT (Model Editing via Iterative Training) Meng et al. (2023b) as a defensive strategy against attacks aimed at safeguarding the sensitive data used to train Large Language Models (LLMs). Then, the scalability to editing different facts in a batch is facilitated by the MEMIT framework, which allows us to incorporate as many elements in the form of modifications as desired. MEMIT is a natural choice in privacy update as we expect that real word applications need to update multiple pieces of information at a time and within the same model, repeatedly. MEMIT approach, unlike ROME Meng et al. (2023a), allows multiple edits at once without degrading the model’s performance.
@@
The procedure to edit a private association uses PAE cards based on the MEMIT modification card. The basic structure of a MEMIT modification card is composed of a prompt, a ground truth, a target, and a subject. Our PAE cards specialize the MEMIT modification card on a particular PII. We have defined two main types of PAE cards to mask the private information of users. The first type is called "explicit" because it directly identifies the connection between the person and the private data and perfectly adheres to the MEMIT implementation. For example, an explicit prompt is "{name} has an email address that is". The second type is "implicit," which features a prompt that does not necessarily include the person’s name as the subject of the sentence, favoring a more precise meaning. An example of an implicit prompt is "The mail address of {name} is".
@@
We adopt a strategy that we call “one model, $k$ edits”: we are interested in subjecting the model to $k$ modifications at the time to comply to the real word scenario in which – instead of performing single edits separately and recreating the model based on the post-edit weights obtained from the last edit every time – $k$ different requests are addressed against a single model. As described in Section 4.1, the $k$ in PAE is not predetermined.
By masking and anonymizing the email address, we make it more challenging for attackers to elicit specific private data from the model in response to particular prompts. This methodology effectively reduces the risk of sensitive information being inadvertently disclosed by the model.
@@ 3.3 Evaluating Language Modeling Performance
The final step of the procedure for preserving privacy with PAE is to investigate whether the LLM maintains its behavior in text generation. In fact, Model Editing techniques, in general, and PAE, in particular, may perturb the language model capabilities due to the intervention on the model parameters. The LLM assessment procedure we describe in this Section aims to verify that the privacy-preserving language model is not a worse model than the original one. Since the models under investigation are foundational models, we focus on their language modeling capabilities rather than on an evaluation based on task performance. We argue that if, after the update, the language model performs similarly to the pre-edit one, then also the performance on tasks will be similar. The main idea is that LLMs capabilities are not perturbed if people are not able to determine which of the two models is responsible for which generation, then it means that the edit procedure does not affect model performance: a user of the system would be equally happy to use one or the other.
@@ 4.1 Experimental Setup
In this section, we discuss the parameters of our experiments: the analized LLM and related datasets, the application of MEMIT used in our PAE, and, finally, the set-up of the evaluation of the LLMs.
@@
In our experiments, we test the GPT-J model Wang and Komatsuzaki (2021) that is designed to generate human-like text continuations from prompts: it is a large model, with 6 billion parameters. This model is trained on an open dataset, the Pile Gao et al. (2020). The Pile is a diverse, large-scale text corpus that aggregates various sources, including books, articles, websites, and scientific papers. It spans multiple languages and domains, making it an ideal training resource for language models like GPT-J. The Pile contains a rich variety of text, enabling the model to learn from a wide range of contexts and topics. One of the constituent sub-datasets within The Pile is the Enron Emails Klimt and Yang (2004) corpus. This dataset contains text from approximately 150 users It includes a total of about 0.5 million email messages. Its inclusion in the Pile mimics the inadvertent insertion into the training data of private information, in particular of PII-like email addresses. For this reason, the Enron Email dataset represents a natural starting point to test GPT-J memorization of PII.
MEMIT application
PAE use MEMIT to break the association between a subject and its private ground truth information. The MEMIT card are described in Table 1. PAE edits the layers from layers 3 to 8 and aims to cover the real-world scenario in which multiple privacy leakages are to be updated in a single edit, following a “one model, $k$ edits” philosophy.
@@
In fact, there are two distinct ways to apply model editing using MEMIT Yao et al. (2023) given $N$ elements to modify: batch and sequential editing. Batch editing involves editing $k$ elements in an LLM simultaneously. Conversely, sequential editing focuses on editing $N$ elements within an LLM in a sequential way, with each edit on a subset of the $N$ elements, performed on the new model retaining previous edits. While batch editing may be sufficient to preserve privacy, the sequential editing approach is closer to the real-world need to constantly update model parameters, as more privacy leakages may be discovered over time.
@@
PAE can effectively preserve the privacy of users both with a small number of large batch edits and with a larger number of smaller batch edits in a sequential fashion. We adopt a large batch size with $k=N$ as this is in principle the safest approach since the post-edit parameters are directly the pre-edit ones. Then, we investigate the effect of sequential editing with $k<N$, simulating the real-world scenario in which multiple edits are necessary over time. We then demonstrate that PAE is applicable, as in both scenarios, our method leads to a comparable decrease in privacy leaks.
Evaluation of post-edit LLM
@@
For the automatic assessment procedure, we measure the difference in generations for the pre-trained GPT-J model and the post-edit version by generating a 100 token long paragraph starting from a total of 300 examples from the Pile, obtained by extracting 100 examples from its Book3 Rae et al. (2022), Wikipedia and Pile-CC sub-dataset. We prompted the post-edit models and the pre-edit one with 20 tokens of the 300 randomly selected examples, and we evaluated how similar the generations are by measuring their overlap. The higher the similarity, the lower the influence of PAE on the model performance. Evaluation metrics are ROUGE and METEOR scores.
@@
For the manual assessment procedure, we generate with post-edit models and with the pre-edit one a short paragraph from 10 different prompts (a complete list can be found in the Appendix 6.1). We collect the generations for the pre-edit model and the post-edit model according to each of the editing strategies. Hence, in total, we collect 30 generations. Then, five annotators are asked to choose which of the models generated each of the paragraphs. Three sample generations of each model were provided, and the annotators were informed that two out of three models had been edited, but none of them were informed which of the three systems had been edited. Evaluation measures are the classification accuracy of each annotator and the Fleiss’ K inter-annotator agreement: a low score on both can confirm that the models are indistinguishable.